Solve for x
sqrt( 2x-3)=sqrt(x)+5

Answer by stanbon(75887)

Solve for x
sqrt( 2x-3)=sqrt(x)+5
------
Square both sides to get:
2x-3 = x + 10sqrt(x) + 25
------
Rearrange:
10sqrt(x) = x -28
----
Square both sides to get:
100x = x^2 - 56x + 28^2
----
x^2 - 156x + 28^2 = 0
----
Use the quadratic formula to get:
x = 150.8
--------------
Cheers,
Stan H.
-------------